OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 18. What is the value of y when x = 4? A.12 B. 16 C. 24 D. 32

OpenStudy (anonymous):

y1/x1^2=y2/x2^2 => 18/9=y/16

OpenStudy (anonymous):

?

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

2=y/16

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so y=16*2=32

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

np